The first-order decomposition of N2O at 1000 K has an initial concentration of 10.9 M. If the concentration after 9.6 s is 7.4 X 10^-3 M, what is the half life?
1
Expert's answer
2019-02-14T04:59:13-0500

For the first-order reaction

log([A]0/[A]) = kt/2.303 and t1/2 = 0.693/k

Find k

log (10.9/7.4*10-3) = k*9.6/2.303

k = 0.76 s-1

Find t1/2

t1/2 = 0.693/0.76 = 0.912 s

Answer: 0.912 s